OpenStudy (anonymous):

which statements are true ?? please help me! √9 is a rational # √-1 is NOT a rational # √ab=√a *√b √a+b=√a+√b 9y√3 and 9√3y are like radicals ^3√5 is the symbol for the square root of 5 √a-b ≠ √a-√b

OpenStudy (anonymous):

You asked this yesterday?!

OpenStudy (anonymous):

and i didnt get help could you help me ?>

OpenStudy (anonymous):

wai!

OpenStudy (anonymous):

Any number that is integer or a fraction can be made by dividing one integer by another is Rational Number. The word comes from "ratio" ( p/q form) The first Option √9 = +3 or -3 as we know (3)^2 = 9 so the answer is integer and we can write it as +3/1 or -3/1 so it is rational!

OpenStudy (anonymous):

√-1 is the complex number i which doesnt have any value and cant be expressed in fraction form So its not rational

OpenStudy (anonymous):

==> √ab=√a *√b check it by taking some arbitrary values for ex by taking a =9 and b = 16 ( took perfect squares for simplification) then √ab = √9*16 = √144 = 12 and √a = √9 = 3 √b =√16 = 4 √a *√b = 3*4 = 12 so what would that makes? √ab=√a *√b is a true relation ( it is an algebra rule that u have to remember!)

OpenStudy (anonymous):

and the same goes for √a+b=√a+√b and √a-b ≠ √a-√b these are also true correct?

OpenStudy (anonymous):

no same does not goes for what u said

OpenStudy (anonymous):

\[\sqrt{a+b}\neq \sqrt{a}+\sqrt{b}\] generally \[\sqrt{a+b}\le \sqrt{a}+\sqrt{b}\]

OpenStudy (anonymous):

Same i told applies for last option! √a-b ≠ √a-√b is true

OpenStudy (anonymous):

Like Radicals : IF the number in the radical and its signs are equal , then we say that the two radicals are like radicals. 9y√3 and 9√3y have same radical sign and number in the radical also is same. therefore these two are like radicals

OpenStudy (anonymous):

square root of 5 is written as ^2√5 not ^3√5 ( ^3√5 represents the cubeth root of 5 )

OpenStudy (anonymous):

Are all clear?

OpenStudy (anonymous):

yes these are all true √9 is a rational # √-1 is NOT a rational # √ab=√a *√b 9y√3 and 9√3y are like radicals √a-b ≠ √a-√b

OpenStudy (anonymous):

you have extremely helpful thank you for taking the time out and explaining it to me